During the hammer throw at a track meet, an 8.0 kg hammer is accidentally thrown straight up. If 784.0 J of work were done on the hammer to give it vertical velocity, how high will it rise?
1
Expert's answer
2013-01-14T11:10:26-0500
The energy conservation law:
T+U = const
T - kinetic energy
U - potential energy
T = m*v^2/2
U = m*g*h
After the throw:
A = T
A + 0 = U + 0
A = m*g*h
h= A/(m*g) = 784/(8*9.8)= 10 m